# Quantifying the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data

### v1

@@ Abstract
Robustness has become a critical attribute for the deployment of RAG systems in real-world applications. Existing research focuses on robustness to explicit noise (e.g., document semantics) but overlooks spurious features (a.k.a. implicit noise). While previous works have explored spurious features in LLMs, they are limited to specific features (e.g., formats) and narrow scenarios (e.g., ICL). In this work, we statistically confirm the presence of spurious features in the RAG paradigm, a robustness problem caused by the sensitivity of LLMs to semantic-agnostic features. Moreover, we provide a comprehensive taxonomy of spurious features and empirically quantify their impact through controlled experiments. Further analysis reveals that not all spurious features are harmful and they can even be beneficial sometimes. Extensive evaluation results across multiple LLMs suggest that spurious features are a widespread and challenging problem in the field of RAG. We release all codes and data at: https://github.com/maybenotime/RAG-SpuriousFeatures.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: An example from the SURE framework (Sec. 3), illustrating the sensitivity of RAG systems to spurious features within grounding data. MD represents the Markdown format. The original retrieved document is fed into the LLMs in different formats, leading to inconsistent responses.
@@
Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm to mitigate LLMs hallucinations (Gao et al., 2023; Yang et al., 2023b), integrating relevant external knowledge to improve the factuality and trustworthiness of LLM-generated outputs (Zhou et al., 2024). However, Retrieval-Augmented Language Models (RALMs) still face substantial robustness issue due to the presence of noise in retrieved documents (Liu et al., 2023; Li et al., 2024c).
@@
Recent research aims to explore the characteristics that affect the robustness of RAG systems from the perspective of grounding data construction (Cuconasu et al., 2024). These studies examine various factors, including the type (Wu et al., 2024a), number (Xu et al., 2024), and position of documents (Liu et al., 2024) within the prompt context. A more detailed discussion of related work is in Appendix A. However, previous analyses primarily focus on explicit noise (i.e., causal features) that significantly alter the semantic information of grounding data(Wu et al., 2024b; Cuconasu et al., 2024), while neglecting spurious features (a.k.a. implicit noise) that make slight modifications but still preserve the meaning of contexts. This limitation extends to existing evaluation benchmarks, which simulate complex noise scenarios to assess the robustness of RAG systems (Chen et al., 2024a; Wu et al., 2024a), yet lack available benchmarks and metrics to measure the robustness of LLMs against spurious features.
@@
Contemporary RAG systems typically employ production-level retrievers, such as Bing and Google, to collect relevant information from the internet. Unlike a single corpus, the internet encompasses diverse data with distinct features. For any given query, there may exist numerous golden documents that contain the correct answer but differ in style, format, or other attributes. As shown in Figure 1, we have observed that LLMs may fail to consistently derive the correct answer from golden documents with different formats. A similar phenomenon is reported in  Sclar et al. (2024) and  He et al. (2024), which demonstrate that LLMs are extremely sensitive to the format of prompts (i.e., spurious features). For more related work, see Appendix A. Unfortunately, there is no statistic and empirical evidence to support the existence of spurious features in the RAG paradigm. This highlights the urgent need to redefine spurious features in RAG and systematically quantify the robustness of RALMs against them.
@@
To address these challenges, we first design a preliminary experiment to demonstrate that RALMs are sensitive to semantic-agnostic features in the grounding data, thereby extending the definition of spurious features to RAG systems. Building on findings from our preliminary experiment and recent studies, we identify five types of spurious features that may appear in RAG scenarios. Then, we propose a novel framework, SURE, for automating the process of robustness evaluation. This framework follows a perturb-then-evaluate approach, offering great scalability. In SURE, automated perturbations are applied on the original instances to inject the corresponding spurious features. The perturbed instances are then examined to ensure that the causal features remain intact. After these steps, we employ tailored metrics to quantify the robustness of RALMs against spurious features. Further analysis reveals that not every spurious features is harmful and they can even be beneficial sometimes. We also find that our proposed metrics are more suitable for evaluating spurious features in golden documents. Based on this finding, we distill the most challenging instances from the synthetic data generated by our framework to create a lighter benchmark, SIG, enabling more efficient robustness evaluation. We evaluate 12 representative LLMs with varying architectures and scales, and the results show that maintaining robustness against spurious features remains a significant challenge, especially for open-source LLMs.
@@
Our contribution can be summarized as follows: 1) We extend the definition of spurious features to RAG systems through a preliminary experiment. To the best of our knowledge, this is the first comprehensive study to define and evaluate spurious features from RAG perspective. 2) We propose a novel evaluation framework, SURE, to assess the robustness of RALMs against spurious features, which effectively simulate five types of spurious features that may appear in real world scenarios. 3) Through extensive experiments and analysis using SURE on two representative LLMs, we provide valuable insights for future research. 4) Based on these findings, we curate a lightweight yet challenging evaluation dataset named SIG and benchmark the robustness of the current state-of-the-art LLMs against spurious features.
2 Preliminary
In this section, we first define causal and spurious features in the context of retrieval-augmented generation and then provide statistical evidence to support the existence of spurious features. Specifically, we start by defining an oracle retriever to sample data according to the preferences of LLMs and then apply statistical testing to validate that RALMs exhibit biases toward semantic-agnostic features within the grounding data.
2.1 Causal and Spurious Features in RAG
In general, causal features are input features that have a direct causal effect on the output of predictive model (Yu et al., 2020). Their relationship is rooted in causality, rather than mere statistical correlation. When it comes to Large Language Models, the meaning and intent of prompts serve as causal features that directly influence the models’ responses. In the context of RAG, causal features refer to the semantic information of grounding data, including whether the correct answer exists, the amount of noise documents in the grounding data, and the content of those noise documents.
In contrast, spurious features are input features that co-occur with causal features and are erroneously captured by the model (Neuhaus et al., 2023). These features exhibit a statistical correlation with the model’s output but lack a causal relationship. Recent research has shown that LLMs are sensitive to seemingly trivial features like prompt formatting, thereby extending the definition of spurious features to LLMs (Sclar et al., 2024). Similarly, we hypothesize that the semantic-agnostic features of the grounding data can be defined as spurious features in RAG systems. However, the conclusion drawn from LLMs may not be applicable to RALMs. Unlike LLMs, the input prompts for RALMs incorporate additional dynamic content—grounding data—to augment their output. This grounding data is derived from retrieved documents rather than static instructions provided directly by users. Therefore, we design a preliminary experiment to validate whether RALMs are sensitive to semantic-agnostic features within the grounding data.
2.2 Oracle Retriever
We aim to confirm the existence of spurious features in RAG scenarios, i.e., to demonstrate the sensitivity of RALMs to semantic-agnostic features in grounding data.
There are some challenges in revealing the sensitivity of RALMs. Specifically, When retrieving from a single corpus, it is difficult to mine semantically equivalent counterparts with obvious differences in semantic-agnostic features. This is because documents from the corpus often share similar styles and formats. With only slight differences in these features, we are difficult to observe significant performance variations using traditional evaluation metrics like accuracy. Thus, more fine-grained metrics are required to capture the subtle performance changes. Inspired by the use of LLMs as supervision signals for document utility (Izacard et al., 2023; Gan et al., 2024), we propose the oracle score, which measures fine-grained performance through calculating the log probability of generating correct answers based on the given documents. The oracle score is defined as follows:
[…]
where $x$ is the input prompt for RALMs, including the instruction $I$, grounding data $G$, and query $Q$; $y$ represents the ground truth answer; $\theta$ denotes the model parameters; and $T$ is the total length of the answer sequence. For cases with multiple answers, we compute the final score by averaging the corresponding oracle scores across all answers. Notably, the oracle score is less effective at estimating the helpfulness of documents when handling queries with long-sequence answers  [^1] In our experiments, we only use queries where the answer sequences is fewer than 5 tokens, ensuring the effectiveness of oracle scores.. This limitation arises from the growing expansion of the solution space as the answer length increases.
We further define an oracle retriever that ranks documents according to their oracle scores. Since the oracle score is computed using LLMs as supervision signals, the oracle retriever can be regarded as a sampler that reflects the LLMs’ preferences for grounding data. With this sampler, we successfully transform the target of demonstrating that RALMs are sensitive to semantic-agnostic features within grounding data into showing that the oracle retriever is biased toward spurious features in documents.
2.3 Preliminary Experiment & Analysis
Using the oracle retriever, we recall 100 documents from the Wikipedia dump for each query in the NQ-open dataset. However, the computational cost of calculating oracle scores across the entire corpus to select top-ranked documents is prohibitive. Therefore, we introduce Contriever-msmarco, a traditional dense retriever, for first-stage retrieval, followed by reranking in descending order based on the oracle scores. In addition to reducing the computational load, this initial retrieval also ensures the semantic similarity of retrieved documents. To further eliminate the effect of causal features, documents without golden answers are filtered out, ensuring that the remaining documents have roughly consistent causal features.
We then select the first-ranked and last-ranked documents for each query from the remaining documents, resulting in two sets, each containing 2658 samples. By comparing the differences in feature distributions between these two sets, we can assess whether the oracle retriever exhibits bias toward semantic-agnostic features. If these two sets do not belong to the same feature distribution, this can be attributed to the oracle retriever’s bias towards semantic-agnostic features during sampling. To confirm that this bias is not introduced by the dense retriever in first-stage retrieval, we establish a control group by randomly sampling two documents instead of selecting the first- and last-ranked documents.
To evaluate whether the two distributions are same, we employ the Kolmogorov-Smirnov (K-S) test. The following semantic-agnostic features are measured in the experiments: 1) Flesh Score, 2) Distinct-1, 3) Dependency Tree Depth, 4) PPL, and 5) Token Length. A detailed introduction of the K-S test and these features can be found in the Appendix B.
We conduct experiments using Mistral-7B-Instruct-v0.3 to implement the oracle retriever. The K-S statistic and P-value are presented in Table 3. Furthermore, we visualize the feature distributions for both the experimental and control groups in Figure 6. For all tested features in the experimental group, the K-S test rejects the null hypothesis, concluding that the distribution of the two sets are significantly different. In contrast, for the control group, the K-S test fails to reject the null hypothesis. The results for Llama-3.1-8B-Instruct are also provided in Appendix B. According to these results, we can conclude that RALMs exhibit bias toward spurious features in documents.
The preliminary experiment provides statistical evidence supporting the existence of spurious features in RAG systems. Nevertheless, it does not offer empirical evidence or quantitative analysis. Inspired by previous data synthesis studies (Tan et al., 2024b; Tong et al., 2024; Li et al., 2024b; Wang et al., 2024a), we use a data synthesis approach to better control feature variables and quantify the robustness of RALMs against spurious features.
@@ 3 Proposed Framework
In this section, we detail our proposed evaluation framework, SURE (Spurious FeatUres Robustness Evaluation), which designed specifically for assessing the robustness of RALMs against spurious features in grounding data. As illustrated in Figure 2, this framework comprise four components: 1) Comprehensive Taxonomy. We identify five distinct types of spurious features that may arise in the context of Retrieval Augmented Generation. 2) Spurious Features Injection. We design a data synthesis pipeline to automate the injection of spurious features, utilizing both model-based and rule-based methods to construct counterparts of the original document with varying spurious features. 3) Causal Features Preservation. We employ a bidirectional entailment algorithm and a string matching strategy to ensure that the causal features of grounding data remain unchanged. 4) Robustness Evaluation. We report the win, lose and robustness rate at instance level to facilitate a fine-grained evaluation.
@@ 3.1 Problem Formulation
Given a query $q$ , the retriever $R$ returns a list of relevant documents from a corpus $D=\{d_{i}\}_{i=1}^{N}$. The relevance between document $d$ and query $q$ can be measured by various methods. In this work, we use a BERT-based dense retriever to obtain the embedding of query and documents, respectively. The relevance score is calculated by computing their dot-product similarity:
$s(q,d_{i})=E(q)\,\raisebox{-0.86108pt}{$\cdot$}\,E(d_{i}).$ (2)
Then, the Top-k documents with the highest similarity scores are retrieved:
@@ 3.2 Taxonomy of Spurious Features
Based on the biased spurious features identified in our preliminary experiment and prior works, we introduce five types of spurious features and their corresponding perturbations in detail as follows.
@@
The same content can be expressed in different styles, using varying tones, words and sentence structures. As shown in Section 2.3, LLMs exhibit biases towards readability-related features. Similarly, for humans, the readability of a text can significantly influence its accessibility to the audience (Yang et al., 2023a). Therefore, we define two perturbations from the perspective of readability style: Simple and Complex. The former simplifies the grounding data by using basic vocabulary and simple sentence structure, while the latter employs professional vocabulary and a formal academic tone to complex the documents.
@@
LLM-generated content, including both misinformation and correct claims, infiltrates every corner of the internet. Recent studies have shown that neural retrievers are biased towards LLM-generated content, leading to the marginalization of human-authored content (Dai et al., 2024; Chen et al., 2024b). Moreover, our preliminary experiments demonstrate that LLMs are biased towards the Perplexity (PPL) of text. Thus, we define two types of source perturbations: LLM-generated and Self-generated. Specifically, the LLM-generated perturbation paraphrases the original document using a powerful LLM, while the self-generated perturbation employs the same backbone model used as the generator in the RAG system.
@@
The arrangement of sentences within a passage typically follows a logical order, ensuring the clarity and coherence in the narrative flow. Here, we simulate scenarios where the intrinsic logical chain is disrupted by three different perturbations: Random, Reverse, and LLM-reranked, each representing a distinct sentence ordering strategy.
@@
The internet contains various data formats, including HTML, Markdown, YAML and JSON. These formats are usually processed into plain text before being fed to LLMs. To mitigate the loss of structural information during this process, some RAG studies propose using the original format, rather than plain text, to augment the generation (Tan et al., 2024a). However, as highlighted in previous research, the prompt format is recognized as a spurious feature that can significantly impact model performance (Sclar et al., 2024; He et al., 2024). Therefore, we perturb the original document with four common formats to explore the impact of grounding data format in the context of RAG.
@@
Metadata is often included in the HTML results returned by search engines. In our framework, we consider two main types of metadata: Timestamp and Data source. The timestamp indicates the time when the data was created, and data source identifies the origin of the data.
@@ 3.3 Spurious Features Injection
An instance is the dynamic component of the prompt $P$, consisting of a query $Q$ and grounding data $G$. The queries are drawn from the NQ-open dataset, while our data source is English Wikipedia dump as of 20 December 2018. To construct the original instances, we first select 1,000 queries based on the close-book QA results of Mistral-7B-Instruct-v0.3 on NQ-open dataset. This subset includes 500 queries that can be answered directly using parametric knowledge (Known) and 500 queries that require external knowledge for answering (Unknown). We then retrieve 100 documents for each query from the Wikipedia dump to serve as the grounding data. Overall, we have a total of 100,000 original instances for the following perturbation step.
@@
As introduced in Section 3.1, the perturbation $g(.)$ injects spurious features by modifying the grounding data. For style and source perturbations, $g(.)$ is implemented using an LLM [^2] Unless otherwise specified, all model-based $g(.)$ are implemented using Llama-3.1-70B-Instruct. prompted by carefully crafted guidelines to modify the raw document, producing counterparts of the original instances. For logic and format perturbations, we develop $g(.)$ as a heuristic method based on a set of predefined rules [^3] One exception is that we implement the LLM-reranked perturbation using an LLM-based $g(.)$.. To simulate metadata that may appear in the real world, we first synthesize pseudo Wikipedia or Twitter links for the raw instances, and then organize them into HTML format using a rule-based $g(.)$. The complete implementation details for automated perturbation are provided in Appendix C.

### v2

# Quantifying and Improving the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data

@@ Abstract
Robustness has become a critical attribute for the deployment of RAG systems in real-world applications. Existing research focuses on robustness to explicit noise (e.g., document semantics) but overlooks implicit noise (spurious features). Moreover, previous studies on spurious features in LLMs are limited to specific types (e.g., formats) and narrow scenarios (e.g., ICL). In this work, we identify and study spurious features in the RAG paradigm, a robustness issue caused by the sensitivity of LLMs to semantic-agnostic features. We then propose a novel framework, SURE, to empirically quantify the robustness of RALMs against spurious features. Beyond providing a comprehensive taxonomy and metrics for evaluation, the framework’s data synthesis pipeline facilitates training-based strategies to improve robustness. Further analysis suggests that spurious features are a widespread and challenging problem in the field of RAG. Our code is available at https://github.com/maybenotime/RAG-SpuriousFeatures.
Quantifying and Improving the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data
Shiping Yang1 2thanks: Work done during internship, Jie Wu2thanks: Correspondence to: jiewu1@microsoft.com, Wenbiao Ding2, Ning Wu2, Shining Liang2, Ming Gong2, Hongzhi Li2, Hengyuan Zhang3, Angel X. Chang1, Dongmei Zhang2 1Simon Fraser University 2Microsoft 3The University of Hong Kong
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm to mitigate LLMs hallucinations Gao et al. (2023); Yang et al. (2023), integrating relevant external knowledge to improve the factuality and trustworthiness of LLM-generated outputs Zhou et al. (2024). However, Retrieval-Augmented Language Models (RALMs) [^1] We focuses on the generation end of RAG systems (i.e., RALMs), and hereafter RAG systems refer to RALMs. still face substantial robustness issue due to the presence of noise in retrieved documents Liu et al. (2023); Li et al. (2024).
Figure 1: An example from the SURE_Wiki dataset (Sec. 4), illustrating the sensitivity of RAG systems to spurious features within grounding data. The original retrieved document is fed into the LLMs in different formats, leading to inconsistent responses.
@@
Recent research aims to explore the characteristics of grounding data that influence the robustness of RAG systems Cuconasu et al. (2024). These studies examine various factors, including the type Wu et al. (2024a), number Xu et al. (2024), and position of documents Liu et al. (2024) within the prompt context. However, previous analyses primarily focus on explicit noise that significantly alter the semantic information (causal features) of grounding data Wu et al. (2024b); Cuconasu et al. (2024), while neglecting implicit noise (spurious features) that introduce semantic-agnostic modifications. This limitation extends to existing evaluation benchmarks, which simulate complex noise scenarios to assess the robustness of RAG systems Chen et al. (2024a); Wu et al. (2024a), yet lack available benchmarks and metrics to measure the robustness of RALMs against spurious features.
@@
Contemporary RAG systems typically employ production-level retrievers, such as Bing and Google, to retrieve relevant information from the internet. Unlike a single corpus, the internet encompasses diverse data with distinct features. For any given query, there may exist numerous golden documents that contain the correct answer but differ in style, format, or other attributes. As shown in Figure 1, we have observed that LLMs may fail to consistently derive the correct answer from golden documents with different formats. A similar phenomenon is reported in  Sclar et al. (2024) and  He et al. (2024), which demonstrate that LLMs are extremely sensitive to the format of prompts (i.e., spurious features). However, spurious features in the RAG paradigm have not been empirically and systematically studied, nor have effective mitigation strategies been proposed.
@@
In this work, we first propose a novel framework, SURE, for automating the process of robustness evaluation. This framework follows a perturb-then-evaluate approach, offering great scalability. In SURE, automated perturbations are applied to the original instances to inject the corresponding spurious features. The perturbed instances are then examined to ensure that the causal features remain intact. After these steps, we employ tailored metrics to quantify the robustness of RALMs against spurious features. To enable more efficient evaluation, we select the most challenging instances to create a lighter benchmark, SIG_Wiki. Extensive evaluations across diverse LLMs and methods indicate that maintaining robustness against spurious features is a widespread and challenging issue. To address this, we introduce two training-based strategies to mitigate the lack of robustness caused by spurious features, leveraging the synthetic data generated by our SURE framework.
@@
Our contribution can be summarized as follows: 1) To the best of our knowledge, this is the first comprehensive study to evaluate spurious features from RAG perspective. We propose a novel evaluation framework, SURE, to assess the robustness of RALMs against spurious features, which includes a comprehensive taxonomy, tailored metrics, and a data synthesis pipeline. 2) Using the synthetic dataset generated by the SURE framework, we curate a lightweight yet challenging evaluation benchmark, SIG_Wiki, and introduce two effective training-based strategies to improve the robustness of RALMs. 3) Further analysis offer valuable insights for future research. For example, we found that not every spurious features is harmful and they can even be beneficial sometimes.
2 Related Work
2.1 Robustness Evaluation of RAG Systems
The retrieved contexts inevitably contains noise in addition to desirable knowledge, which may mislead LLMs to produce an incorrect response Bian et al. (2024); Feldman et al. (2024). Previous works have explored automated evaluation frameworks to assess the robustness of RAG systems in various settings Chen et al. (2024a); Cuconasu et al. (2024). Wu et al. (2024a) provided a detailed taxonomy of noise documents to further simulate the complexity of real-world scenarios and highlighted the potential positive effects of certain types of noise.
While these studies have identified several explicit noises that affect the robustness of RAG systems, they predominantly overlook implicit noises. Even some works evaluate the influence of implicit noise, they are often limited to specific types, such as typos Cho et al. (2024) and formatting Tan et al. (2024a). In this work, we comprehensively study semantic-agnostic noises (i.e, spurious features) in RAG systems.
2.2 Prompt Sensitivity of LLMs
Prompts are instructions provided to an LLM to perform specific tasks automatically and ensure desired qualities in the generated output. However, it is known that current LLMs are sensitive to the features of input prompts Zhu et al. (2023). This sensitivity poses challenges for researchers attempting to evaluate the model’s performance accurately and precisely Zhuo et al. (2024). Beyond causal features that significantly influence the meaning of prompts Wei et al. (2022); Kong et al. (2024), existing works have demonstrated that LLMs are highly sensitive to spurious features Sclar et al. (2024) in non-RAG scenarios, e.g, prompt formatting He et al. (2024), language style Li et al. (2023), the order of options Pezeshkpour and Hruschka (2024).
In contrast, our work focuses on spurious features within the context of RAG. Unlike non-RAG scenarios (e.g., ICL), where spurious features appear in user-provided instructions and remain static across queries, in RAG they occur in retrieved documents and vary dynamically for each query.
@@ 3 Proposed Framework
In this section, we detail our proposed evaluation framework, SURE (Spurious FeatUres Robustness Evaluation), which designed specifically for assessing the robustness of RALMs against spurious features in grounding data. As illustrated in Figure 2, this framework comprise four components: 1) Comprehensive Taxonomy. We identify and define five common types of spurious features in RAG scenarios. 2) Spurious Features Injection. We design a data synthesis pipeline to automate the injection of spurious features, utilizing both model-based and rule-based methods to construct counterparts of the original document with varying spurious features. 3) Causal Features Preservation. We employ a bidirectional entailment algorithm and a string matching strategy to ensure that the causal features of grounding data remain unchanged. 4) Robustness Evaluation. We introduce three metrics (Win Rate, Lose Rate, and Robustness Rate) to facilitate fine-grained, instance-level evaluation.
@@ 3.1 Problem Formulation
Given a query $q$ , the retriever $R$ returns a list of relevant documents from a corpus $D=\{d_{i}\}_{i=1}^{N}$. The relevance between document $d$ and query $q$ can be measured by various methods. In this work, we use a BERT-based dense retriever to obtain the embedding of query and documents, respectively. The relevance score is calculated by computing their dot-product similarity. Then, the Top-k documents with the highest similarity scores are retrieved:
@@ 3.2 Taxonomy of Spurious Features
We develop a comprehensive taxonomy of spurious features, informed by our preliminary experiments (see Appendix A) and insights from prior research. The five types of spurious features and their corresponding perturbations are detailed below.
@@
The same content can be expressed in different styles, using varying tones, words and sentence structures. As shown in Appendix A.2, LLMs exhibit biases towards readability-related features. Therefore, we define two perturbations from the perspective of readability style: Simple and Complex. The former simplifies the grounding data by using basic vocabulary and simple sentence structure, while the latter employs professional vocabulary and a formal academic tone to complex the documents.
@@
Recent studies have shown that neural retrievers are biased towards LLM-generated content, leading to the marginalization of human-authored content Dai et al. (2024); Chen et al. (2024b). Moreover, our preliminary experiments demonstrate that LLMs are biased towards the Perplexity (PPL) of text. Thus, we define two types of source perturbations: LLM-generated and Self-generated. Specifically, the LLM-generated perturbation paraphrases the original document using a powerful LLM, while the self-generated perturbation employs the same backbone model used as the generator in the RAG system.
@@
In RAG systems, documents are often segmented into multiple chunks and may be retrieved in varying orders. Here, we simulate scenarios where the intrinsic logical flaw is disrupted by three different perturbations: Random, Reverse, and LLM-reranked, each representing a distinct sentence ordering strategy.
@@
The internet contains various data formats, including HTML, Markdown, YAML and JSON. These formats are usually processed into plain text before being fed to LLMs. To mitigate the loss of structural information during this process, some RAG studies propose using the original format, rather than plain text, to augment the generation Tan et al. (2024a). However, as highlighted in previous research, the prompt format is recognized as a spurious feature that can significantly impact model performance Sclar et al. (2024); He et al. (2024). Therefore, we perturb the original document with four common formats to explore the impact of grounding data format.
@@
Metadata is often included in the HTML results returned by search engines. In our framework, we focus on two types: Timestamp and Data source. The timestamp marks when the data was created, and the data source indicates its origin  [^2] Timestamp may serve as causal features in time-sensitive tasks. In our experiments, the NQ dataset generally does not contain time-sensitive queries.. For timestamp perturbations, pre and post denote whether the timestamp is before or after the LLM’s knowledge cutoff date. For data source perturbations, wiki and twitter represent the domains of the URLs.
@@ 3.3 Spurious Features Injection
An instance is the dynamic component of the prompt $P$, consisting of a query $Q$ and grounding data $G$. To construct the original instances, we first select 1,000 queries from the NQ-open dataset. For each query, we then retrieve 100 documents from the Wikipedia dump to serve as grounding data, yielding 100,000 instances for the following perturbation step.
@@
As introduced in Section 3.1, the perturbation $g(.)$ injects spurious features by modifying the grounding data. For style and source perturbations, $g(.)$ is implemented using an LLM [^3] Unless otherwise specified, all model-based $g(.)$ are implemented using Llama-3.1-70B-Instruct. prompted by carefully crafted guidelines to modify the raw document, producing counterparts of the original instances. For logic and format perturbations, we develop $g(.)$ as a heuristic method based on a set of predefined rules. To simulate real-world metadata, we first synthesize pseudo Wikipedia or Twitter links for the raw instances, and then organize them into HTML format using a rule-based $g(.)$. Further details for automated perturbation are provided in Appendix B.